OpenStudy (anonymous):

Marie has 9 novels and 4 biographies. 12 mysteries and 14 magazines. What fractional part of the reading material is either novels or magazines?

OpenStudy (anonymous):

add all the materials = 9 + 4 + 12 +14 = 29 if novels = 9/29 if magazines = 14/29

OpenStudy (anonymous):

then you add the 14 and 9 and whatever the answer is you get __ over 29 right?

OpenStudy (anonymous):

?